OpenStudy (anonymous):

A cyclist travels 40 km at a speed of x km/h. Find the time taken in terms of x. Find the time taken when his speed is reduced by 2 km/h. If the difference between the times is 1 hour. Find the original speed x.

OpenStudy (ajprincess):

\[speed=\frac{ distance }{ time }\] \[x=\frac{40}{time}\] \[time=\frac{40}{x}\] After the speed is reduced by 2km/h the new speed will be x-2km/h \[speed=\frac{ distance }{ time }\] \[x-2=\frac{40}{time}\] \[time_1=\frac{40}{x-2}\] \[time-time_1=1\] \[\frac{40}{x}-\frac{40}{x-2}=1\] Can u solve for x nw?

OpenStudy (anonymous):

x=10

OpenStudy (ajprincess):

Yup u r right. The last eqn I gav u to solve shud be \[\frac{40}{x-2}-\frac{40}{x}=1\] and when u solve this u will get x=10